Please walk me throught this: (Imaguinary solutions of quadratic equations)
Solve: 2x^2 + 1 = 0

Found 2 solutions by stanbon, bluerose22:
Answer by stanbon(75887)

Please walk me throught this: (Imaginary solutions of quadratic equations)
Solve: 2x^2 + 1 = 0
--------------------------
2x^2 = -1
x^2 = -1/2
x = +/- [sqrt(2)/2)i
=========================
Cheers,
Stan H.

Answer by bluerose22(2)

2x^2+1=0
4x+1=0
-1 -1
4x=-1
1x=-0.25
x=-.25